OpenStudy (anonymous):

Prove that (1+3ω+ω^2) (1+ω+2ω^2)=2

hartnn (hartnn):

whats 1+w^2 from 1+w+w^2=0 ??

OpenStudy (anonymous):

i think we have to multiply all

hartnn (hartnn):

but it can be simplified without multiplication.....

OpenStudy (anonymous):

how

hartnn (hartnn):

(1+3w+w^2)= ((1+w^2)+3w) now from 1+w+w^2=0,we have (1+w^2)=-w so, ((1+w^2)+3w) = (-w+3w) = 2w got this? can u do the same for (1+ω+2ω^2)

OpenStudy (anonymous):

(1+3ω+ω^2)(1+ω+2ω^2) arranging it: (1+ω^2+3ω)(1+ω+2ω^2) putting formulas: ((-ω)+3ω) ((-ω^2)+2ω^2)

OpenStudy (anonymous):

Simplify the terms in the bracket..

hartnn (hartnn):

u are on right track,go on...

OpenStudy (anonymous):

*brackets

OpenStudy (anonymous):

(2ω)(ω^2)

OpenStudy (anonymous):

You can further simplify it, I believe.

OpenStudy (anonymous):

2ω^3=2

OpenStudy (anonymous):

You did it ...

OpenStudy (anonymous):

if theres a problem in solving further questions i'll ask it so be there thanks